If U = I + 3j - 2k and V = 4i - 2j - 4k are vectors, find 3U + V
1
Expert's answer
2013-09-19T07:59:49-0400
If U = I + 3j - 2k and V = 4i - 2j - 4k are vectors, find 3U + V
Answer: 3u+v=3(i + 3j - 2k)+4i - 2j - 4k=7i+7j-10k